countryboy:

The spherical container has a radius of 21 cm. What is the approximate volume of the container? Use 3.14 to approximate pi. Round to the nearest hundredth if necessary. Enter your answer in the box. cm³

Shadow:

\[V_{sphere} = \frac{ 4 }{ 3 } \pi r^{3}\]

Shadow:

Do you know how to use this formula?

countryboy:

no

Shadow:

r = radius input 21 in for r, then solve

Shadow:

\[V = \frac{ 4 }{ 3 } \pi (21)^{3}\]

Shadow:

\[V = \frac{ 4 }{ 3 } \pi (9261)\] \[V = 12,348 \pi = 38,792.39\]

Shadow:

Don't forget cm^3